OpenStudy (ria23):

A lake is stocked with 2,300 young trout. After x years, the function P(x) = 2,300 e^-0.2x provides the number of the original trout that are still in the lake. When will there be 500 of the original trout left? I got stuck... ;-;

OpenStudy (ria23):

\[P(x)=2,300e^{-0.2x}\]\[500=2300e^{-0.2x}\] \[\frac{ 2300 }{ 500 }=4.6\]

OpenStudy (anonymous):

should be 500/2300 = e^(-0.2x)

OpenStudy (anonymous):

ln(5/23) = -0.2x x = ln(5/23)/(-0.2) x = 7.63 years

OpenStudy (ria23):

Thank yhu... How did yhu get 5/23?

OpenStudy (anonymous):

reduced 500/2300

OpenStudy (ria23):

Oh. Well, thank yhu for helping n_n